### v1

@@ Abstract
This paper investigates regularized stochastic gradient descent (SGD) algorithms for estimating nonlinear operators from a Polish space to a separable Hilbert space. We assume that the regression operator lies in a vector-valued reproducing kernel Hilbert space induced by an operator-valued kernel. Two significant settings are considered: an online setting with polynomially decaying step sizes and regularization parameters, and a finite-horizon setting with constant step sizes and regularization parameters. We introduce regularity conditions on the structure and smoothness of the target operator and the input random variables. Under these conditions, we provide a dimension-free convergence analysis for the prediction and estimation errors, deriving both expectation and high-probability error bounds. Our analysis demonstrates that these convergence rates are nearly optimal. Furthermore, we present a new technique for deriving bounds with high probability for general SGD schemes, which also ensures almost-sure convergence. Finally, we discuss potential extensions to more general operator-valued kernels and the encoder-decoder framework.
@@ 1 Introduction
In this paper, we consider a general model abstracted from nonlinear operator learning problems:
$y=h^{\dagger}(x)+\epsilon.$ (1.1)
Here, $\mathcal{X}$ is a Polish space and $\left(\mathcal{Y},\langle\cdot,\cdot\rangle_{\mathcal{Y}},\|\cdot\|_{\mathcal{Y}}\right)$ is a separable Hilbert space. The pair $(x,y)$ satisfying (1.1) is a random variable taking values in $\mathcal{X}\times\mathcal{Y}$, distributed according to an unknown probability measure $\rho$. The operator $h^{\dagger}:\mathcal{X}\rightarrow\mathcal{Y}$ is a measurable (possibly nonlinear) mapping defined by the conditional expectation $h^{\dagger}(x):=\mathbb{E}[y|x]$. The noise term $\epsilon$ is a centered $\mathcal{Y}$-valued random variable, assumed to be independent of $x$ and to have finite variance, i.e., $\sigma^{2}:=\mathbb{E}[\|\epsilon\|^{2}_{\mathcal{Y}}]<\infty$.
The model (1.1) has been widely employed in surrogate approaches for structured output prediction [46, 21, 26, 7, 14, 4]. In practice, many applications involve inputs or outputs with explicit or implicit discrete structures. Examples of implicitly structured data used in predictions include text, images, and videos in document processing and retrieval, as well as genes and proteins in computational biology. To learn models that predict outputs with structured components, surrogate methods embed the structured output into a Hilbert space, which formulates the task as operator regression with an infinite-dimensional output space. During prediction, a decoding step maps the output from the Hilbert space back to the original structured output space. Structured prediction tasks such as image completion [46], label ranking [29], and graph prediction [6] can thus be addressed through operator learning using surrogate approaches. Another important application of model (1.1) is functional output regression [24, 28, 25, 41]. These problems have become increasingly relevant with the growing capacity to collect functional data, motivating a shift toward a functional perspective in modeling [25]. This has led to the development of the now-thriving field of operator learning [1, 30], which aims to approximate operators between Hilbert (or more generally, Banach) spaces using data. A prominent example is the learning of solution operators for partial differential equations (PDEs), where the goal is to approximate the operator that maps a parameter space—describing the physical and geometrical constraints of the PDE—to its solution space [33, 3, 36]. In this paper, we study operator learning algorithms designed to act directly on functions rather than on high-dimensional vectors. This functional perspective allows us to capture the intrinsic properties of the problem, avoiding reliance on specific discretizations or pixelizations.
@@
This simple construction of operator-valued kernels has been adopted in previous works, e.g., [4, 1]. Note that all elements $h$ in the vector-valued RKHS $\mathcal{H}$ are measurable. We also consider a more general class of kernels in Section 3.1. In particular, our analysis covers most operator-valued kernels, including those of the form (1.2) with a compact operator $W$. Furthermore, when choosing kernels as in (1.2), it follows from [12, Example 5] that $K$ is a Mercer [resp. $\mathcal{C}_{0}$] [^2] That is, the Banach space of continuous functions vanishing at infinity with the uniform norm. kernel if $\mathcal{K}$ is Mercer [resp. $\mathcal{C}_{0}$], implying that all operators in $\mathcal{H}$ are continuous.
@@ 2 Main Results
The proof of Proposition 2.1 is deferred to Appendix A.1. By applying Proposition 2.1 with $W=I$—in which case the kernel coincides with that specified by Assumption 1—the iteration (1.5) can be equivalently expressed as
@@ 2.1 Assumptions
A sufficient (though not necessary) condition for this assumption is that $u_{k}=O(k^{-\frac{1}{s}-\epsilon})$ for some $\epsilon>0$. For a detailed discussion on this condition, we refer the reader to [22]. In this paper, Assumption 3 is introduced to derive sharper error bounds. When combined with Assumption 3 for some $10<s<$ 1, it leads to improved convergence rates. This condition—commonly known as the capacity condition—was first introduced in [17] and has since been widely adopted in the literature, including [38, 4, 22, 41], as a way to capture the intrinsic complexity of infinite-dimensional learning problems. Assumptions 2 and 3 are essential to establish dimension-free convergence analysis. As we will show, the resulting convergence rates depend explicitly on the parameters $r$ and $s$, reflecting the regularity of the target operator and the capacity of the input random variables, respectively.
@@
According to [41, Proposition 2.1], this assumption is equivalent to
@@ 2.2 Error Bounds in Expectation
Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 4 are satisfied. Define $\{h_{t}\}_{t\geq 1}$ through (2.1) with step sizes $\{\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}\}_{t\geq 1}$ and regularization parameters $\{\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}\}_{t\geq 1}$, where $10<\theta_{1}<$ 1, $10<\theta_{2}<$ 1 and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$. Additionally, let $t_{0}$ satisfy $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, $t_{0}\geq\exp\{\frac{1}{\theta_{1}}\}$, and
@@
$\mathbb{E}_{z^{T}}\left[\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})\right]\leq c_{1,1}\begin{cases}(T+t_{0})^{-\theta_{1}},&\text{when }s<1,\\ (T+t_{0})^{-\theta_{1}}\log(T+t_{0}),&\text{when }s=1.\\ \end{cases}$
@@
In Theorem 2.3, since the constant $c_{4}$ is independent of $t_{0}$, one can choose $t_{0}$ sufficiently large to satisfy the required conditions. Compared to Assumption 3 with $s=1$, the stronger assumption with $10<s<$ 1 only removes a logarithmic factor in the convergence rate. It is also clear that the convergence rate saturates at $r=1/2$, i.e., increasing $r$ beyond $1/2$ does not yield further improvement. According to [41, Theorem 2.9], the result is minimax optimal (up to a logarithmic term) when $s=1$ and $r<1/2$. Compared to the unregularized SGD algorithm analyzed in [41, Theorem 2.4], adding a regularization term here leads to faster convergence. Specifically, while the prediction error rate of unregularized SGD in [41] saturates at $r=(1-s)/2$, the regularized SGD in our work improves the saturation level to $r=1/2$.
@@
The convergence rate of the estimation error saturates at $r=1$, which improves the convergence of unregularized SGD in [41, Theorem 2.6], where the rate saturates at $r=\frac{1-s}{2}$. Moreover, in [22, Theorem 3] and [41, Theorem 2.6], we cannot guarantee the convergence of the estimation error with decaying step sizes for $s=1$, whereas adding a regularization term addresses this issue. According to [41, Theorem 2.9], the convergence rate with decaying step sizes is minimax optimal when $r<1$.
@@ 2.3 High-probability Error Bounds
$\mathcal{E}(h_{T+1})-\mathcal{E}(h^{\dagger})$ $\leq c_{2,3}\begin{cases}T^{-\theta_{3}}\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{2}T\log^{4}\frac{2}{\delta},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T\log^{2}\frac{2}{\delta}+T^{1-3\theta_{3}}\log^{3}T\log^{4}\frac{2}{\delta},&\text{ when }s=1,\\ \end{cases}$
$\lesssim\log^{4}\frac{2}{\delta}\begin{cases}T^{-\theta_{3}},&\text{ when }s<1,\\ T^{-\theta_{3}}\log T,&\text{ when }s=1.\end{cases}$
@@ 3.1 Extension to General Kernel Setting
    Let $\mathcal{H}$ be separable, which is true if the spaces $\mathcal{X}$ and $\mathcal{Y}$ are separable and $K$ is a Mercer kernel [11, Corollary 5.2]. A kernel $K$ is Mercer if and only if the RKHS induced by $K$ is a subspace of the spaces of continuous operators from $\mathcal{X}$ to $\mathcal{Y}$, which in turn holds if and only if $K$ is locally bounded and $K(x,\cdot)$ is strongly continuous for any $x\in\mathcal{X}$ [11, Proposition 5.1].
@@
    We further assume that $K$ is strongly measurable. Under assumptions (1) and (2), this is equivalent to requiring that each element in $\mathcal{H}$ is a measurable function [11, Proposition 3.3].
@@ 3.2 Application to Structured Prediction
for any estimator $h$, as detailed in [13, 4]. Let $\hat{f}=D(\hat{h})$, denote the estimator for $f^{\dagger}$ obtained via the algorithm. The surrogate approach for structured prediction is illustrated in Figure 1.
@@ 3.3 Combining with PCA Encoder-decoder Framework
The primary function of PCA is to extract the principal features of the data. High-dimensional data often suffers from the curse of dimensionality, and dimensionality reduction—achieved by identifying and retaining the most significant information—serves as an effective remedy. This constitutes the central role of PCA. In our setting, we employ PCA to reduce the samples from an infinite-dimensional space to a finite-dimensional one. The PCA algorithm applied to a random input $x$ in $\mathcal{X}$ seeks to minimize the reconstruction error $\mathbb{E}\left[\left\|(I-P)x\right\|_{\mathcal{X}}^{2}\right]$ over $\Pi_{d_{\mathcal{X}}}$, the set of all orthogonal projections $P$ with rank $d_{\mathcal{X}}$. Given the covariance operator of $x$ defined as $\Sigma_{x}:=\mathbb{E}[x\otimes x]$, there exist eigenvalue-eigenvector pairs $\{\lambda_{i}^{d_{\mathcal{X}}},\phi_{i}^{d_{\mathcal{X}}}\}_{i\geq 1}$ satisfying $\langle\Sigma_{x}\phi_{i}^{d_{\mathcal{X}}},\phi_{j}^{d_{\mathcal{X}}}\rangle_{\mathcal{X}}=\delta_{ij}$ and $\lambda_{1}^{d_{\mathcal{X}}}\geq\lambda_{2}^{d_{\mathcal{X}}}\geq\cdots\geq 0$, where $\delta_{ij}=1$ if $i=j$, otherwise 00. It can be shown that the optimal PCA projection is given by
@@
It then follows that $\mathbb{E}\left[\left\|(I-P_{d_{\mathcal{X}}}^{\mathcal{X}})x\right\|_{\mathcal{X}}^{2}\right]=\sum_{i>d_{\mathcal{X}}}\lambda_{i}^{d_{\mathcal{X}}}$, see [32, Theorem 3.8].
In practice, it is usually difficult to obtain $\Sigma_{x}$ directly, so we typically use the empirical covariance operator $\Sigma_{x}^{T}=\frac{1}{T}\sum_{i=1}^{T}x_{i}\otimes x_{i}$ as a substitute, thus deriving the empirical PCA. Following the same procedure, we naturally obtain the empirical encoder ^$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. Similarly, by replacing $\Sigma_{x}^{T}$, the input random variable $x$, and the rank $d_{\mathcal{X}}$ with $\Sigma_{y}^{T}:=\frac{1}{T}\sum_{i=1}^{T}y_{i}\otimes y_{i}$, the output random variable $y$ and $d_{\mathcal{Y}}$, respectively, we apply empirical PCA to $y$ in $\mathcal{Y}$ with rank $d_{\mathcal{Y}}$, and obtain the empirical encoder ^$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$.
$\mathcal{X}$$\mathcal{Y}$$\mathbb{R}^{d_{\mathcal{X}}}$$\mathbb{R}^{d_{\mathcal{Y}}}$${\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}h^{\dagger}}$^$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$^$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$f^{\dagger}$
Figure 2: Commutative diagram of PCA enceoder-decoder framework
We now formulate the estimator as $h:=\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ f\circ\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. It is then natural to choose $f^{\dagger}=\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ h^{\dagger}\circ\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. This formulation naturally give rise to a commutative diagram, as illustrated in Figure 2. The works [3, 31] represent $f$ using neural networks. In constract, we represent $f$ in an RKHS induced by a matrix-valued kernel $k:\mathbb{R}^{d_{\mathcal{X}}}\times\mathbb{R}^{d_{\mathcal{X}}}\to\mathcal{B}(\mathbb{R}^{d_{\mathcal{Y}}})$. Specifically, we consider kernels of the form $k(u,v)=\phi(\|u-v\|_{\mathbb{R}^{d_{\mathcal{X}}}})$, where $\phi:[0,\infty)\to\mathbb{R}$ is a radial function, such that $k$ is positive definite for $d_{\mathcal{X}}>0$. This property can be equivalently characterized by requiring $\phi$ to be completely monotone. Notable examples satisfying this condition include the inverse multiquadrics $\phi(x)=\left(c^{2}+x^{2}\right)^{-\beta}$ and the Gaussian kernel $\phi(x)=e^{-\alpha|x|^{2}}$ for any $c>0$, $\beta>0$, and $\alpha>0$; see [45]. Let $\mathcal{H}_{k}$ denote the RKHS induced by the matrix-valued kernel $k$. With a slight abuse of notation, we define the prediction error as
@@
where $\eta$ is the step size. This can be interpreted as an SGD scheme based on the samples {^$\{\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x_{t},\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}y_{t}\}_{t=1}^{T}$ in $\mathcal{H}_{k}$. Accordingly, we define $h_{1}=0$ and

### v3

@@ Abstract
We consider a class of statistical inverse problems involving the estimation of a regression operator from a Polish space to a separable Hilbert space, where the target lies in a vector-valued reproducing kernel Hilbert space induced by an operator-valued kernel. To address the associated ill-posedness, we analyze regularized stochastic gradient descent (SGD) algorithms in both online and finite-horizon settings. The former uses polynomially decaying step sizes and regularization parameters, while the latter adopts fixed values. Under suitable structural and distributional assumptions, we establish dimension-independent bounds for prediction and estimation errors. The resulting convergence rates are near-optimal in expectation, and we also derive high-probability estimates that imply almost sure convergence. Our analysis introduces a general technique for obtaining high-probability guarantees in infinite-dimensional settings. Possible extensions to broader kernel classes and encoder-decoder structures are briefly discussed.
@@ 1 Introduction
In this work, we consider the nonlinear regression model
$y=h^{\dagger}(x)+\epsilon,$ (1.1)
where $x$ belongs to a Polish space $\mathcal{X}$, and $y$ lies in a separable Hilbert space $\mathcal{Y}$. The pair $(x,y)$ is distributed according to an unknown probability measure $\rho$ on $\mathcal{X}\times\mathcal{Y}$. The regression operator $h^{\dagger}:\mathcal{X}\to\mathcal{Y}$ is defined as the conditional expectation
$h^{\dagger}(x)=\mathbb{E}[y|x],$
and the noise $\epsilon$ is a centered $\mathcal{Y}$-valued random variable with finite variance $\sigma^{2}:=\mathbb{E}[\|\epsilon\|^{2}_{\mathcal{Y}}]$, assumed independent of $x$. The objective is to recover $h^{\dagger}$ from independent, identically distributed samples of $(x,y)$.
This formulation encompasses a broad class of statistical inverse problems in infinite-dimensional settings, where the forward map $h^{\dagger}$ must be recovered from noisy, indirect observations. Traditional approaches to inverse problems often employ explicit regularization techniques, such as Tikhonov regularization or spectral filtering, to stabilize the solution, typically under certain assumptions on the forward operator’s structure. In our work, we also use explicit regularization by penalizing the norm in a vector-valued reproducing kernel Hilbert space (RKHS). However, unlike classical deterministic frameworks, our approach studies stochastic gradient descent algorithms in infinite-dimensional function spaces with streaming data, requiring new analytical tools to establish convergence and error bounds under operator regularity conditions.
A key motivation for this model arises from structured output prediction, where the outputs exhibit functional or combinatorial structure. In many practical scenarios, the output space is not finite-dimensional but consists of complex objects such as functions, sequences, or graphs. To address these tasks, surrogate learning approaches embed the structured output into a Hilbert space and learn a continuous operator mapping inputs to embedded outputs. Applications include image completion [43], label ranking [27], graph prediction [6], and protein function prediction [14]. This framework also includes functional output regression, which has become increasingly relevant with the growing availability of functional data in science and engineering. Here, both the input and output may be infinite-dimensional, motivating the development of operator-theoretic regression approaches. These applications have spurred the growth of the operator learning paradigm, which aims to learn mappings between function spaces directly from data. A prominent example is learning solution operators for parameterized partial differential equations (PDEs), where the operator maps coefficient or boundary condition functions to PDE solutions [30, 3, 33]. In this paper, we focus on learning algorithms that operate directly on function spaces rather than on finite-dimensional discretizations. This functional viewpoint preserves the intrinsic geometry and infinite-dimensional structure of the problem, enabling discretization-invariant analysis and broad applicability. Our setting naturally leads to the study of regularized learning in vector-valued RKHSs, where the regression operator is modeled within a rich nonparametric hypothesis space induced by operator-valued kernels.
@@
This simple construction of operator-valued kernels has been adopted in previous works, e.g., [4, 1]. Note that all elements $h$ in the vector-valued RKHS $\mathcal{H}$ are measurable. We also consider a more general class of kernels in Section 3.1. In particular, our analysis covers most operator-valued kernels, including those of the form (1.2) with a compact operator $W$. Furthermore, when choosing kernels as in (1.2), it follows from [11, Example 5] that $K$ is a Mercer [resp. $\mathcal{C}_{0}$] [^2] That is, the Banach space of continuous functions vanishing at infinity with the uniform norm. kernel if $\mathcal{K}$ is Mercer [resp. $\mathcal{C}_{0}$], implying that all operators in $\mathcal{H}$ are continuous.
@@ 2 Main Results
The proof of Proposition 2.1 is deferred to Appendix A.1. By applying Proposition 2.1 with W=IW=I—in which case the kernel coincides with that specified by Assumption 1—the iteration (1.5) can be equivalently expressed as
@@ 2.1 Assumptions
A sufficient (though not necessary) condition for this assumption is that $u_{k}=O(k^{-\frac{1}{s}-\epsilon})$ for some $\epsilon>0$. For a detailed discussion on this condition, we refer the reader to [21]. In this paper, Assumption 3 is introduced to derive sharper error bounds. When combined with Assumption 3 for some $0<s<1$, it leads to improved convergence rates. This condition—commonly known as the capacity condition—was first introduced in [17] and has since been widely adopted in the literature, including [35, 4, 21, 38], as a way to capture the intrinsic complexity of infinite-dimensional learning problems. Assumptions 2 and 3 are essential to establish dimension-free convergence analysis. As we will show, the resulting convergence rates depend explicitly on the parameters $r$ and $s$, reflecting the regularity of the target operator and the capacity of the input random variables, respectively.
@@
According to [38, Proposition 2.1], this assumption is equivalent to
@@ 2.2 Error Bounds in Expectation
Suppose that Assumption 1, Assumption 2, Assumption 3 and Assumption 4 are satisfied. Define $\{h_{t}\}_{t\geq 1}$ through (2.1) with step sizes $\{\eta_{t}=\bar{\eta}(t+t_{0})^{-\theta_{1}}\}_{t\geq 1}$ and regularization parameters $\{\lambda_{t}=\bar{\lambda}(t+t_{0})^{-\theta_{2}}\}_{t\geq 1}$, where $0<\theta_{1}<1$, $0<\theta_{2}<1$ and $\bar{\eta}\bar{\lambda}>\theta_{2}\min\{r,1\}$. Additionally, let $t_{0}$ satisfy $(t_{0}+1)^{\theta_{1}}\geq\bar{\eta}(\kappa^{2}+\bar{\lambda})$, $t_{0}\geq\exp\{\frac{1}{\theta_{1}}\}$, and
@@
[…]
@@
In Theorem 2.3, since the constant $c_{4}$ is independent of $t_{0}$, one can choose $t_{0}$ sufficiently large to satisfy the required conditions. Compared to Assumption 3 with $s=1$, the stronger assumption with $0<s<1$ only removes a logarithmic factor in the convergence rate. It is also clear that the convergence rate saturates at $r=1/2$, i.e., increasing $r$ beyond $1/2$ does not yield further improvement. According to [38, Theorem 2.9], the result is minimax optimal (up to a logarithmic term) when $s=1$ and $r<1/2$. Compared to the unregularized SGD algorithm analyzed in [38, Theorem 2.4], adding a regularization term here leads to faster convergence. Specifically, while the prediction error rate of unregularized SGD in [38] saturates at r=(1s)/2r=(1-s)/2, the regularized SGD in our work improves the saturation level to $r=1/2$.
@@
The convergence rate of the estimation error saturates at $r=1$, which improves the convergence of unregularized SGD in [38, Theorem 2.6], where the rate saturates at $r=\frac{1-s}{2}$. Moreover, in [21, Theorem 3] and [38, Theorem 2.6], we cannot guarantee the convergence of the estimation error with decaying step sizes for $s=1$, whereas adding a regularization term addresses this issue. According to [38, Theorem 2.9], the convergence rate with decaying step sizes is minimax optimal when $r<1$.
@@ 2.3 High-probability Error Bounds
[…]
@@ 3.1 Extension to General Kernel Setting
    Let $\mathcal{H}$ be separable, which is true if the spaces $\mathcal{X}$ and $\mathcal{Y}$ are separable and $K$ is a Mercer kernel [12, Corollary 5.2]. A kernel $K$ is Mercer if and only if the RKHS induced by $K$ is a subspace of the spaces of continuous operators from $\mathcal{X}$ to $\mathcal{Y}$, which in turn holds if and only if $K$ is locally bounded and $K(x,\cdot)$ is strongly continuous for any $x\in\mathcal{X}$ [12, Proposition 5.1].
@@
    We further assume that $K$ is strongly measurable. Under assumptions (1) and (2), this is equivalent to requiring that each element in $\mathcal{H}$ is a measurable function [12, Proposition 3.3].
@@ 3.2 Application to Structured Prediction
[…]
@@ 3.3 Combining with PCA Encoder-decoder Framework
The primary function of PCA is to extract the principal features of the data. High-dimensional data often suffers from the curse of dimensionality, and dimensionality reduction—achieved by identifying and retaining the most significant information—serves as an effective remedy. This constitutes the central role of PCA. In our setting, we employ PCA to reduce the samples from an infinite-dimensional space to a finite-dimensional one. The PCA algorithm applied to a random input $x$ in $\mathcal{X}$ seeks to minimize the reconstruction error $\mathbb{E}\left[\left\|(I-P)x\right\|_{\mathcal{X}}^{2}\right]$ over $\Pi_{d_{\mathcal{X}}}$, the set of all orthogonal projections $P$ with rank $d_{\mathcal{X}}$. Given the covariance operator of $x$ defined as $\Sigma_{x}:=\mathbb{E}[x\otimes x]$, there exist eigenvalue-eigenvector pairs $\{\lambda_{i}^{d_{\mathcal{X}}},\phi_{i}^{d_{\mathcal{X}}}\}_{i\geq 1}$ satisfying $\langle\Sigma_{x}\phi_{i}^{d_{\mathcal{X}}},\phi_{j}^{d_{\mathcal{X}}}\rangle_{\mathcal{X}}=\delta_{ij}$ and $\lambda_{1}^{d_{\mathcal{X}}}\geq\lambda_{2}^{d_{\mathcal{X}}}\geq\cdots\geq 0$, where $\delta_{ij}=1$ if $i=j$, otherwise 0. It can be shown that the optimal PCA projection is given by
@@
It then follows that $\mathbb{E}\left[\left\|(I-P_{d_{\mathcal{X}}}^{\mathcal{X}})x\right\|_{\mathcal{X}}^{2}\right]=\sum_{i>d_{\mathcal{X}}}\lambda_{i}^{d_{\mathcal{X}}}$, see [28, Theorem 3.8].
In practice, it is usually difficult to obtain $\Sigma_{x}$ directly, so we typically use the empirical covariance operator $\Sigma_{x}^{T}=\frac{1}{T}\sum_{i=1}^{T}x_{i}\otimes x_{i}$ as a substitute, thus deriving the empirical PCA. Following the same procedure, we naturally obtain the empirical encoder $\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. Similarly, by replacing $\Sigma_{x}^{T}$, the input random variable $x$, and the rank $d_{\mathcal{X}}$ with $\Sigma_{y}^{T}:=\frac{1}{T}\sum_{i=1}^{T}y_{i}\otimes y_{i}$, the output random variable $y$ and $d_{\mathcal{Y}}$, respectively, we apply empirical PCA to $y$ in $\mathcal{Y}$ with rank $d_{\mathcal{Y}}$, and obtain the empirical encoder $\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$ and empirical decoder $\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$.
$\mathcal{X}$$\mathcal{Y}$$\mathbb{R}^{d_{\mathcal{X}}}$$\mathbb{R}^{d_{\mathcal{Y}}}$$\color[rgb]{0.82,0.01,0.11}\definecolor[named]{pgfstrokecolor}{rgb}{0.82,0.01,0.11}h^{\dagger}$$\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$$\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}$$f^{\dagger}$
Figure 2: Commutative diagram of PCA encoder-decoder framework
We now formulate the estimator as $h:=\widehat{\mathcal{D}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ f\circ\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. It is then natural to choose $f^{\dagger}=\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}\circ h^{\dagger}\circ\widehat{\mathcal{D}}_{d_{\mathcal{X}}}^{\mathcal{X}}$. This formulation naturally gives rise to a commutative diagram, as illustrated in Figure 2. The works [3, 29] represent $f$ using neural networks. In contrast, we represent $f$ in an RKHS induced by a matrix-valued kernel $k:\mathbb{R}^{d_{\mathcal{X}}}\times\mathbb{R}^{d_{\mathcal{X}}}\to\mathcal{B}(\mathbb{R}^{d_{\mathcal{Y}}})$. Specifically, we consider kernels of the form $k(u,v)=\phi(\|u-v\|_{\mathbb{R}^{d_{\mathcal{X}}}})$, where $\phi:[0,\infty)\to\mathbb{R}$ is a radial function, such that $k$ is positive definite for $d_{\mathcal{X}}>0$. This property can be equivalently characterized by requiring $\phi$ to be completely monotone. Notable examples satisfying this condition include the inverse multiquadrics $\phi(x)=\left(c^{2}+x^{2}\right)^{-\beta}$ and the Gaussian kernel $\phi(x)=e^{-\alpha|x|^{2}}$ for any $c>0$, $\beta>0$, and $\alpha>0$; see [42]. Let $\mathcal{H}_{k}$ denote the RKHS induced by the matrix-valued kernel $k$. With a slight abuse of notation, we define the prediction error as
@@
where $\eta$ is the step size. This can be interpreted as an SGD scheme based on the samples $\{\widehat{\mathcal{E}}_{d_{\mathcal{X}}}^{\mathcal{X}}x_{t},\widehat{\mathcal{E}}_{d_{\mathcal{Y}}}^{\mathcal{Y}}y_{t}\}_{t=1}^{T}$ in $\mathcal{H}_{k}$. Accordingly, we define $h_{1}=0$ and